how many kilograms of water must be added to 10 kg of a 75% antifreeze solution in order to produce a 50% solution?
Answer by stanbon(57940)

how many kilograms of water must be added to 10 kg of a 75% antifreeze solution in order to produce a 50% solution?
----------------------------------
Let "x" be the amount of water added.
Equation:
alcohol + alcohol = alcohol
0.75(10) + 0(x) = 0.50(10+x)
Multiply thru by 100 to get:
75*10 = 50*10 + 50x
50x = 25*10
x = 5 kg (amount of water needed in the mixture)
====================================================
Cheers,
Stan H.